OpenStudy (anonymous):

find the y-intercept of the line passing through the point (5,-6) with a slope of -1/7

OpenStudy (anonymous):

i recommend using point slope form to solve this do you know what that is?

OpenStudy (anonymous):

\[y-y _{1}=m(x-x _{1})\]

OpenStudy (anonymous):

where y1 = y coordinate, m = slope, x1 = x coordinate

OpenStudy (anonymous):

plug in the information you are given into that and solve for y. not y1, just y.

OpenStudy (anonymous):

after you have done that your equation will be in y=mx + or - b where b is your y - intercept

OpenStudy (anonymous):

can you just give the answer please @TechnoSoul

OpenStudy (anonymous):

y + 6 = -1/7(x -5) y+6 = -1/7x + 5/7 y= -1/7x - 37/7

OpenStudy (anonymous):

-37/7 or -5 2/7 is your y- intercept

OpenStudy (anonymous):

thank you so much you very smart